OpenStudy (anonymous):

2 question test. will medal and fan. Please help.

OpenStudy (xmissalycatx):

??

OpenStudy (anonymous):

i know. it is short but very important. half my grade.

OpenStudy (xmissalycatx):

Help with what..

OpenStudy (nysha):

what are the questions

OpenStudy (anonymous):

math and 1. Which example supports the given conjecture? The sum of two even numbers is an even number. A. 1 + 3 = 4 B. 4 + 6 = 10 C. 4 + 5 = 9 D. 0 + 1 = 1

pooja195 (pooja195):

Hi :) Welcome to openstudy Im sorry but we don't help with test questions.

OpenStudy (xmissalycatx):

Even numbers are 0 ,2, 4, 6, 8 Odd numbers are 1,3, 5, 7, 9

OpenStudy (anonymous):

oh, ok. thank you anyway.

OpenStudy (nysha):

sorry

OpenStudy (nysha):

yes